### v1

@@ Abstract
This paper develops theory and algorithms for interacting large language model agents (LLMAs) using methods from statistical signal processing and microeconomics. While both fields are mature, their application to decision-making by interacting LLMAs remains unexplored. Motivated by Bayesian sentiment analysis on online platforms, we construct interpretable models and stochastic control algorithms that enable LLMAs to interact and perform Bayesian inference. Because interacting LLMAs learn from both prior decisions and external inputs, they can exhibit bias and herding behavior. Thus, developing interpretable models and stochastic control algorithms is essential to understand and mitigate these behaviors.
@@ I Introduction
This paper develops theory and algorithms for interacting Large Language Model Agents (LLMAs) by leveraging techniques from signal processing and microeconomics. Specifically, we focus on developing interpretable models and stochastic control algorithms for LLMAs, enabling them to interact sequentially for Bayesian inference.
We construct interpretable models of LLMAs at two levels of abstraction, as outlined in Figure 1. First, we model an individual LLMA as a rationally inattentive Bayesian utility maximizer, capturing the agent’s decision-making process under limited attention. Second, we extend this approach to a sequence of LLMAs engaging in Bayesian social learning, where each agent acts as a Bayesian utility maximizer. Our models are inspired by the self-attention mechanism in large language models (LLMs) and observed challenges, such as model collapse, that can arise during LLM training.
@@
Figure 1: Summary of the proposed contributions: We discuss the different black-box models for LLMA and how LLMAs can be used as a sensing mechanism to perform Bayesian inference. Part 1 models the LLMAs as a rationally inattentive Bayesian utility maximizer and numerically establishes the behavior in applications of product quality identification and hate speech classification. Part 2 discusses how Bayesian social learning in a sequence of LLMAs can be used for sequential state estimation. However, in Part 3, we show that the agents can perform the same incorrect action due to herding. We then discuss a stochastic control approach to delay herding when LLMAs are centrally controlled and when they are autonomous but are incentivized.
Furthermore, motivated by the observed bias in the behavior of interacting LLMAs, we demonstrate that a sequence of LLMAs engaging in Bayesian social learning converge to identical estimates, or ”herd”. To address this phenomenon, we propose a stochastic control approach, formulating an optimal stopping problem to balance the trade-off between privacy and herding, to detect the failure state. Our approach is designed for two scenarios: (a) when the LLMAs are centrally controlled, and (b) when they operate autonomously.
Our goal is to demonstrate that concepts from controlled sensing and microeconomics, traditionally applied to human decision-making, can be used to both understand and synthesize the behavior of interacting LLMAs [12, 7, 13, 39, 40]. We support our theoretical findings with numerical experiments using advanced LLMs for Bayesian inference on real-world data. This paper is crafted to engage a broad readership, highlighting applications of Bayesian agents in diverse fields, including financial news analysis, e-commerce review evaluation, and online toxicity detection. These examples underscore the flexibility of our methodologies for cross-disciplinary applications. The reproducible code for our experiments is publicly accessible at github.com/aditj/sociallearningllm.
@@ I-A Motivation
LLM agents (LLMAs) are being rapidly deployed for different applications and to quote Sam Altman, CEO of OpenAI (creators of ChatGPT, a popular LLM which has 200 million weekly active users): “2025 is when (AI) agents will work”.
LLMAs use a large language model (LLM) to parse the input and have additional agency to perform tasks. LLMs (such as ChatGPT and Llama) are neural networks with billions of parameters trained on trillions of tokens of textual data to parse long texts for summarizing, compiling key facts, and generating new text. The key technical improvement that leads to the efficient deployment of LLMs is the transformer architecture [75]. The effectiveness of LLMs on textual texts has made their deployment and adoption widespread [54]. Many applications have been proposed in healthcare, online platform moderation, and finance, where these LLMs are used to parse the textual observations and suggest decisions based on their outputs [48].
@@
Figure 2: Engineering with large language model agents: We propose engineer with LLMs on three different levels: a) First, we propose constructing an LLMA with an LLM attached to the Bayesian engine. The LLM acts as a sensor for the text input and outputs interpretable low-dimensional outputs, which are used by the Bayesian engine to produce a state estimate. b) We formulate the necessary and sufficient conditions for a LLMAs to be a rationally inattentive Bayesian utility maximizer (RIBUM). We also present algorithms to reconstruct feasible utilities and rational inattention costs if the LLMA is indeed a RIBUM, attributing the LLMA with an interpretable microeconomic model. c) We show how a sequence of LLMAs can efficiently perform sequential Bayesian social learning by controlling their outputs to delay herding optimally. Our Bayesian social learning models can be extended to study Bayesian social learning in a network of LLMAs.
@@ I-A1 Interacting large language model agents
It is predicted that by 2025, 90% of web content will be generated by large language models (LLMs) [1]. In recent practical implementations, individual LLMs are part of a bigger system, referred to as LLMAs, and interact with the content generated by other LLMAs and the external environment [86]. Furthermore, recent research has shown how generative models are trained on the data generated by other generative models can collapse [67]. Therefore naturally, LLMAs interact with each other either implicitly or explicitly.
@@ I-A2 Interpretable Engineering of LLMAs
Therefore, it becomes imperative to study interpretable models for these agents since many of the proposed applications these agents involve sensitive information (like personal records, financial information, bio-medical data, and personal preferences). By interpretable, we refer to models that facilitate a transparent understanding of complex models through clear and explainable representations of their decision-making processes. The workflows of the AI agents also include making decisions, and the interpretability and reliability of these agents become vital for them to be trustworthy. Therefore, mathematical models are needed to aid in engineering and deploying LLMAs. To this end, we propose a LLMA composed of an LLM and a Bayesian engine, which by construction is interpretable. Further, we use Bayesian revealed preferences  [^1] The framework of Bayesian revealed preferences is also referred to as inverse optimization or inverse reinforcement learning. to reconstruct a Bayesian utility function for both our constructed LLMA and for off-the-shelf LLMAs.
@@ I-B Main Results
Sec. 3: LLM Agent as a SensorSec. 4: LLM Agent as a Rationally Inattentive Bayesian Utility Maximizer (RIBUM)Part 1Indiviual LLM Agent Part 2Network of LLM Agents Sec. 5: Sequential Bayesian Social LearningSec. 6: Word-of-MouthBayesian Social LearningSec. 6: AsynchronousBayesian Social LearningPart 3Stochastic Control for delaying Herding in LLM Agents Sec. 7: Optimal Stopping for Centrally Controlled LLM AgentsSec. 8: Optimal Stopping for Incentivized Autonomous LLMAsSec. 9: Stochastic Approximation for Optimal Policy Sec. 10: Numerical Results on Bayesian inference Product Quality and Hate Speech Peddler Identification
@@
This paper builds on tools from Bayesian revealed preferences from microeconomics (inverse reinforcement learning), sequential Bayesian estimation (from signal processing), and structured stochastic control (from control theory) to construct interpretable models and synthesize interaction of LLMAs. The impact of our results on more efficient, systematic, and interpretable engineering of LLMAs is summarized in Figure 2. The main contributions of this paper are:
@@
    We propose constructing a LLMA as a composition of a large language model (LLM) sensor, which acts a low-dimensional map from the text space and a Bayesian engine, which uses the measurement from the LLM to update the posterior and act optimally. We show how this model is useful for interpretable Bayesian inference with applications in sequential data on online platforms.
@@
    To obtain an interpretable utility function for a LLMA, we provide necessary and sufficient conditions in Theorem 1 for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). For a LLMA who is a RIBUM, we propose Algorithm 2 and Algorithm 3 to reconstruct the max-margin and sparsest utility estimate, respectively. Our methods are applicable both our LLMA and off-the-shelf LLMAs.
@@
    To delay herding in a sequence of LLMAs, we formulate an optimal stopping problem for two regimes: a) when the LLMAs are centrally controlled by an entity b) when the LLMAs are autonomous but are incentivizes by an entity. We show in Theorem 3 and Theorem 4 that under certain assumptions on the observation matrix and cost functions, the optimal policy for the partially observed Markov decision process of both the optimal stopping problems has a threshold structure. We then propose a stochastic approximation algorithm in Algorithm 7, which exploits the structural results to estimate the optimal policy parameters. The algorithm does not need access to the system parameters, is computationally efficient, and can track changes in the system.
@@
    We finally present several numerical experiments to demonstrate the efficacy of our proposed methods. We show how our constructed LLMA can be used for interpretable Bayesian inference for analyzing financial data. We show how the Bayesian revealed preferences framework can estimate the utility of an off-the-shelf LLM when used for hate-speech detection. Finally we show numerical studies on two examples of sequential Bayesian inference: hate speech peddler identification and product quality analysis, to demonstrate herding of LLMAs, and applicability of our structural results.
@@
    What is a principled approach to analyze whether a LLMA is a Bayesian utility maximizer and also reconstructs its utility function given only black-box access?
@@
    How does one systematically study Bayesian social learning in multiple interacting LLMAs to explain observed behaviours such as herding and model collapse?
@@ I-C Organization
This paper is organized into three parts, and the schematic of the organization is given in Figure 3. Part I discusses interpretable model for a single LLMA and attempt at answering questions 1 and 2 above. Section II discusses the related work in large language models, agents using LLMs, and current interpretable models for Bayesian inference. Section III discusses the mathematical model used for modeling LLMAs in this paper and motivates the different components involved. Section IV gives the necessary and sufficient conditions for the LLMAs to be rationally inattentive Bayesian utility maximizers (RIBUM). It furthers proposes algorithms to estimate the utility function for a LLMA which is a RIBUM.
@@
To decrease the bias when a sequence of LLMAs perform Bayesian inference, Part III deals with stochastic control for delaying herding in interacting LLMAs performing Bayesian sequential learning proves structural results, and proposes a stochastic approximation approach. Section IX considers the problem of a central controller optimally optimizing a sequence of autonomous LLMAs to achieve the state estimation by optimally controlling herding. Section X proposes a stochastic approximation based approach to approximate the optimal policy, which has a threshold switching curve. Numerical results on real-life text classification tasks and related applications are discussed in Section XI. Section XII concludes the paper with discussions on future works, open problems, and research opportunities. The appendix contains the proofs and details about the numerical experiments.
For the ease of the reader, we have included a motivation and a discussion subsection in each section, which grounds the different aspects of LLMAs to a real-life application and different microeconomics and statistical signal processing tools presented in the section. We also provide different block diagrams and illustrative examples to further aid the reader.
@@ II Brief Literature Review and Related Work
In this section, we review some of the substantial related work on LLMs [^2] In this paper, LLMs also refer to the various transformer architectures that process multi-modal data, including images, audio, and documents., LLMAs and social learning using LLMs. We first provide a brief background on LLMs and discuss the different applications and models for LLMAs. We provide motivation for the interpretability of the LLM agents. Finally, we review literature in sequential state estimation setup studied in this paper and provide motivation for using a sequence of contemporary LLMAs in a classical Bayesian inference setting. We also review applications of sequential Bayesian inference using LLMAs. Table I summarizes some of the related work.
@@ II-A Background on LLMs
Large language models (LLMs) have become omnipresent in various industry applications in different aspects, given the drastic improvement in compute availability and rapid development of open and closed-sourced models [63]. They are being rapidly deployed for various applications, including education, information retrieval, gaming, recommendation systems, and understanding graphs [36, 64, 34, 21].
The primary reason for their proliferation is that they are able to ingest a chunk of multi-modal data and provide useful inferences from them. This is possible because they are deep learning networks (transformer architecture) with billions of parameters trained on massive amounts (in the order of petabytes) of data (CommonCrawl, etc.) using extremely fast GPUs that can parallelize computations efficiently. There are two different classes of LLMs: a) open source LLMs like Llama and Mistral [33, 74] and b) closed source LLMs like ChatGPT and Claude. Open source LLMs make available the underlying deep learning architecture that they use, some even share the data that the LLM is trained on; closed-source LLMs on the other hand only provide an inferencing interface, where the LLM can be asked different questions, the answer to which is provided by the LLM. The methods and framework of our paper only require black-box access to these LLMs and are generally applicable to both classes.
@@ II-B LLM Agents
Standalone LLMs are powerful tools for many applications, but recent work has proposed using LLMs as part of bigger systems and integrating them into existing workflows. These LLMs often have agency and are referred to as LLM agents (LLMAs) [24]. There are two distinct features that LLMAs have that make them different from LLMs:
@@
    Decisions: In the workflows that the LLMAs are used in, they are provided agency using different mechanisms, including function calls

### v2

@@ Abstract
This paper discusses the theory and algorithms for interacting large language model agents (LLMAs) using methods from statistical signal processing and microeconomics. While both fields are mature, their application to decision-making involving interacting LLMAs remains unexplored. Motivated by Bayesian sentiment analysis on online platforms, we construct interpretable models and stochastic control algorithms that enable LLMAs to interact and perform Bayesian inference. Because interacting LLMAs learn from both prior decisions and external inputs, they can exhibit bias and herding behavior. Thus, developing interpretable models and stochastic control algorithms is essential to understand and mitigate these behaviors.
@@ I Introduction
This paper discusses the theory and algorithms for interacting Large Language Model Agents (LLMAs) by leveraging techniques from Bayesian inference, stochastic control, and microeconomics. Specifically, we focus on developing interpretable models and stochastic control algorithms for LLMAs, enabling them to interact sequentially for Bayesian inference.
We construct interpretable models of LLMAs at two levels of abstraction, as outlined in Figure 1. First, we model an individual LLMA as a rationally inattentive Bayesian utility maximizer, capturing the agent’s decision-making process under limited attention. Second, we extend this approach to a sequence of LLMAs engaging in Bayesian social learning or groupthink, where each agent is a Bayesian utility maximizer. Our models are inspired by the self-attention mechanism in large language models (LLMs) and observed challenges, such as model collapse, that arise during LLM training.
@@
Figure 1: Summary of the proposed contributions: We discuss the different blackbox models for LLMA and how LLMAs can be used as a sensing mechanism to perform Bayesian inference. Part 1 models the LLMAs as a rationally inattentive Bayesian utility maximizer and numerically establishes the behavior in applications of product quality identification and hate speech classification. Part 2 discusses how Bayesian social learning in a sequence of LLMAs can be used for sequential state estimation. However, in Part 3, we show that the agents can perform the same incorrect action due to herding. We then discuss a stochastic control approach to delay herding when LLMAs are centrally controlled and when they are autonomous but are incentivized.
Furthermore, motivated by the observed bias in the behavior of interacting LLMAs, we demonstrate that a sequence of LLMAs engaging in Bayesian social learning end up making identical decisions, or ”herd”. To address this phenomenon, we propose a stochastic control approach, formulating an optimal stopping problem to balance the trade-off between privacy and herding, to detect the failure state. Our approach is designed for two scenarios: (a) when the LLMAs are centrally controlled, and (b) when they operate autonomously.
The overarching goal of this paper is to demonstrate that concepts from controlled sensing and microeconomics, traditionally applied to human decision-making, can be used to both understand and synthesize the behavior of interacting LLMAs [1, 2, 3, 4, 5]. We support our theoretical findings with numerical experiments using advanced LLMs for Bayesian inference [6] on real-world data. This paper is written to engage a broad readership, highlighting applications of Bayesian agents in diverse fields, including financial news analysis, e-commerce review evaluation, and online toxicity detection. These examples underscore the flexibility of our methodologies for cross-disciplinary applications. The reproducible code for our experiments is publicly accessible at github.com/aditj/sociallearningllm.
@@ I-A Motivation
LLM agents (LLMAs) are being rapidly deployed for different applications [7], and to quote Sam Altman, CEO of OpenAI (creators of ChatGPT, a popular LLM which has 200 million weekly active users): “in 2025, we may see the first AI agents join the workforce” [8].
LLMAs use a large language model (LLM) to parse the input and have additional functionality to perform tasks. LLMs (such as ChatGPT and LLaMA) are neural networks with billions of parameters trained on trillions of tokens of textual data to parse long texts for summarizing, compiling key facts, and generating new text. The key technical improvement that leads to the efficient deployment of LLMs is the transformer architecture [9]. The effectiveness of LLMs on textual texts has made their deployment and adoption widespread [10]. Many applications have been proposed in healthcare, online platform moderation, and finance, where these LLMs are used to parse the textual observations and suggest decisions based on their outputs [11].
@@
Figure 2: Engineering with large language model agents (LLMAs): We propose engineering with LLMs on three different levels: a) First, we construct LLMAs with an LLM attached to the Bayesian engine. The LLM acts as a sensor for the text input and outputs interpretable low-dimensional outputs, which are used by the Bayesian engine to produce a state estimate. b) Second, we formulate necessary and sufficient conditions for a LLMAs to be a rationally inattentive Bayesian utility maximizer (RIBUM). We also present algorithms to reconstruct feasible utilities and rational inattention costs if the LLMA is indeed a RIBUM, attributing the LLMA with an interpretable microeconomic model. c) Finally, we demonstrate how a sequence of LLMAs can efficiently perform sequential Bayesian social learning by controlling their outputs to delay herding optimally. Our Bayesian social learning models can be extended to study Bayesian social learning in a network of LLMAs.
@@ I-A1 Interacting large language model agents
It is predicted that by 2030, 90% of web content will be generated by large language models (LLMs) [13]. In recent practical implementations, individual LLMs are part of a bigger system, referred to as LLMAs, and interact with the content generated by other LLMAs and the external environment [14]. Furthermore, recent research has shown how generative models are trained on the data generated by other generative models can collapse [15]. Therefore, naturally, LLMAs interact with each other either implicitly or explicitly.
@@ I-A2 Interpretable Engineering of LLMAs
Therefore, it becomes imperative to study interpretable models for these agents since many of the proposed applications for these agents involve sensitive information (like personal records, financial information, bio-medical data, and personal preferences). By interpretable, we refer to models that facilitate a transparent understanding of complex models through clear and explainable representations of their decision-making processes. The workflows of the AI agents also include making decisions, and the interpretability and reliability of these agents become vital for them to be trustworthy. Therefore, mathematical models are needed to aid in engineering and deploying LLMAs. To this end, we propose a LLMA composed of an LLM and a Bayesian engine, which by construction is interpretable. Further, we use Bayesian revealed preferences [^1] The framework of Bayesian revealed preferences is also referred to as inverse optimization or inverse reinforcement learning. to reconstruct a Bayesian utility function for both our constructed LLMA and for off-the-shelf LLMAs.
@@ I-B Main Results
Sec. 3: LLM Agent as a SensorSec. 4: LLM Agent as a Rationally Inattentive Bayesian Utility Maximizer (RIBUM)Part 1Individual LLM Agent Part 2Network of LLM Agents Sec. 5: Sequential Bayesian Social LearningSec. 6: Word-of-MouthBayesian Social LearningSec. 6: AsynchronousBayesian Social LearningPart 3Stochastic Control for delaying Herding in LLM Agents Sec. 7: Optimal Stopping for Centrally Controlled LLM AgentsSec. 8: Optimal Stopping for Incentivized Autonomous LLMAsSec. 9: Stochastic Approximation for Optimal Policy Sec. 10: Numerical Results on Bayesian inference Product Quality and Hate Speech Peddler Identification
@@
This paper builds on Bayesian revealed preferences from microeconomics (inverse reinforcement learning), sequential Bayesian estimation (from signal processing), and structured stochastic control to construct interpretable models and synthesize interaction of LLMAs. The impact of our results on more efficient, systematic, and interpretable engineering of LLMAs is summarized in Figure 2. The main contributions of this paper are:
@@
    We propose constructing a LLMA as a composition of a large language model (LLM) sensor [^2] We consider LLMs as social sensors and not physical sensors since they do not sense physical quantities like temperature but virtually analyze text and multimodal data to provide observations., which acts as a low-dimensional map from the text space, and a Bayesian engine, which uses the measurement from the LLM to update the posterior and act optimally. We show how this model is useful for interpretable Bayesian inference with applications in sequential data on online platforms.
@@
    To obtain an interpretable utility function for a LLMA, we provide necessary and sufficient conditions in Theorem 1 for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). For a LLMA who is a RIBUM, we propose Algorithm 2 and Algorithm 3 to reconstruct the max-margin and sparsest utility estimate, respectively. Our methods are applicable to both our LLMA and off-the-shelf LLMAs.
@@
    To delay herding in a sequence of LLMAs, we formulate an optimal stopping problem for two regimes:
    a) when the LLMAs are centrally controlled by an entity b) when the LLMAs are autonomous but are incentivized by an entity. We show in Theorem 3 and Theorem 4 that under certain assumptions on the observation matrix and cost functions, the optimal policy for the partially observed Markov decision process of both the optimal stopping problems has a threshold structure. We then propose a policy gradient algorithm in Algorithm 7, which exploits the structural results to estimate the optimal policy parameters. The algorithm does not need access to the system parameters, is computationally efficient, and can track changes in the system.
@@
    We finally present several numerical experiments to demonstrate the efficacy of our proposed methods. We show how our constructed LLMA can be used for interpretable Bayesian inference for analyzing financial data. We show how the Bayesian revealed preferences framework can estimate the utility of an off-the-shelf LLM when used for hate-speech detection. Finally, we show numerical studies on two examples of sequential Bayesian inference: hate speech peddler identification and product quality analysis, to demonstrate herding of LLMAs, and the applicability of our structural results.
@@
    What is a principled approach to analyze whether a LLMA is a Bayesian utility maximizer and also reconstructs its utility function given only blackbox access?
@@
    How does one systematically study Bayesian social learning in multiple interacting LLMAs to explain observed behaviors such as herding and model collapse?
@@ I-C Organization
This paper is organized into three parts, and the schematic of the organization is given in Figure 3. Part I discusses the interpretable model for a single LLMA and attempts to answer questions 1 and 2 above. Section II discusses the related work in large language models, agents using LLMs, and current interpretable models for Bayesian inference. Section III discusses the mathematical model used for modeling LLMAs in this paper and motivates the different components involved. Section IV gives the necessary and sufficient conditions for the LLMAs to be rationally inattentive Bayesian utility maximizers (RIBUM). It further proposes algorithms to estimate the utility function for a LLMA which is a RIBUM.
@@
To decrease the bias when a sequence of LLMAs perform Bayesian inference, Part III deals with stochastic control for delaying herding in interacting LLMAs performing Bayesian sequential learning proves structural results, and proposes a policy gradient approach. Section IX considers the problem of a central controller optimally optimizing a sequence of autonomous LLMAs to achieve the state estimation by optimally controlling herding. Section X proposes a policy gradient based approach to approximate the optimal policy, which has a threshold switching curve. Numerical results on real-life text classification tasks and related applications are discussed in Section XI. Section XII concludes the paper with discussions on future works, open problems, and research opportunities. The appendix contains the proofs and details about the numerical experiments.
We emphasize that the paper is built around a coherent framework with the unifying theme of building interpretable models for interacting LLMAs using Bayesian social learning and powerful generative models from behavioral economics. For the ease of the reader, we have included a motivation and a discussion subsection in each section, which grounds the different aspects of LLMAs to a real-life application and different microeconomics and statistical signal processing tools presented in the section. We also provide different block diagrams and illustrative examples to further aid the reader. All the results reported in this paper are fully reproducible with code downloadable from github.com/aditj/sociallearningllm.
@@ II Brief Literature Review and Related Work
In this section, we review related work on LLMs [^3] In this paper, LLMs also refer to the various transformer architectures that process multi-modal data, including images, audio, and documents., LLMAs and social learning using LLMs. We first provide a brief background on LLMs and discuss the different applications and models for LLMAs. We provide motivation for the interpretability of the LLM agents. Finally, we review literature in sequential state estimation setup studied in this paper and provide motivation for using a sequence of contemporary LLMAs in a classical Bayesian inference setting. We also review applications of sequential Bayesian inference using LLMAs. Table I summarizes some of the related work.
@@ II-A Background on LLMs
Large language models (LLMs) have become omnipresent in various industry applications, given the drastic improvement in compute availability and rapid development of open and closed-sourced models [20]. They are being rapidly deployed for various applications, including education, information retrieval, gaming, recommendation systems, and understanding graphs [23, 24, 21, 22].
The primary reason for the proliferation of LLMs is that they can take a high-dimensional multi-modal (text, images, audio etc.) data input and provide useful inferences about them. This is possible because they are deep learning networks (transformer architecture) with billions of parameters trained on massive amounts (in the order of petabytes) of data (CommonCrawl, etc.) using extremely fast GPUs that can parallelize computations efficiently. There are two different classes of LLMs: a) open source LLMs like LLaMA and Mistral [57, 58] and b) closed source LLMs like ChatGPT and Claude. Open source LLMs make available the underlying deep learning architecture that they use, some even share the data that the LLM is trained on; closed-source LLMs on the other hand only provide an inferencing interface, where the LLM can be asked different questions, the answer to which is provided by the LLM. The methods and framework of our paper only require blackbox access to these LLMs and are generally applicable to both classes.
@@ II-B LLM Agents
Standalone LLMs are powerful tools for many applications, but there is strong motivation to consider networked LLMs as part of complex systems and integrate them into existing workflows. These networked LLMs are often given functionality to interact with the virtual environment and are referred to as LLM agents (LLMAs) [25]. There are two distinct features that LLMAs have that make them different from LLMs:
@@
    Decisions: In the workflows that the LLMAs are used in, they are provided functionality (agency) using different mechanisms, including function calls